A football is kicked in the air and lands on the ground 32m from where it was kicked. If the maximum height reached by the ball is 14m, determine the transformational form of the quadratic function representing the path of the football while it is in the air.
Answer by jrfrunner(365)

You need to put the formula in standard form y=a(x-h)^2+k
(h,k) is the vertex
-
you have three points. the origion (0,0) the end (32,0) and its middle (16,14) which is the vertex
-
let's use input the vertex in the standard formula
y=a(x-16)^2+14
-
lets use the origin (0,0) to solve for "a"
0=a(-16)^2+14
0=256a+14
a=-14/256= - 7/128
--
substitute a into formula
y=-7/128*(x-16)^2+14